OpenStudy (anonymous):

. Is a triangle with sides of length 6 ft., 21 ft., 23 ft. a right triangle? Explain

OpenStudy (anonymous):

use Pythagorean theorem

OpenStudy (anonymous):

Just plug the numbers into the pythagorean theorem, if the two numbers equal each other, then yes, if not, no. a^2 + b^2 = c^2 where a = 6, b = 21, and c = 23 so 6^2 + 21^2 = 23^2 36 + 441 = 529 447 does not equal 529 No, it is not a right triangle